# RAG-DDR: Optimizing Retrieval-Augmented Generation Using Differentiable Data Rewards

### v1

@@ 4 Experimental Methodology
This section first describes datasets, evaluation metrics, and baselines. Then we introduce the implementation details of our experiments. More experimental details are shown in Appendix A.1.
@@
Dataset. In our experiments, we follow RA-DIT (Lin et al., 2023) and use the instruction tuning datasets for training and evaluating RAG models. For all datasets and all baselines, we use bge-large (Xiao et al., 2023) to retrieve documents from the MS MARCO 2.0 (Bajaj et al., 2016).
@@
During the training of DDR, we collect ten datasets covering two tasks, open-domain QA and reasoning. Specifically, we randomly sample 32,805 samples for the training set and 2,000 samples for the development set in our experiments. Following previous work (Lin et al., 2023; Xu et al., 2024a), we select the knowledge-intensive tasks for evaluation, including open-domain question answering, multi-hop question answering, slot filling, and dialogue tasks. The open-domain QA tasks consist of NQ (Kwiatkowski et al., 2019), MARCO QA (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017), which require models to retrieve factual knowledge to help answer the given question. For more complex tasks, such as multi-hop QA and dialogue, we use HotpotQA dataset (Yang et al., 2018) and Wikipedia of Wizard (WoW) (Dinan et al., 2019) for evaluation. Besides, we also employ T-REx (Elsahar et al., 2018) to measure one-hop fact look-up abilities of models.
@@
Evaluation. Following Xu et al. (2024a), we utilize Rouge-L and F1 as evaluation metrics for MARCO QA task and WoW task, respectively. For the rest tasks, we use Accuracy.
Baselines. In our experiments, we compare DDR with five baseline models, including zero-shot models and supervised finetuning models.
@@
We first treat the LLM as a black box and conduct three baselines, including LLM w/o RAG, Vanilla RAG and REPLUG. For the LLM w/o RAG model, we directly feed the query to the LLM and ask it to produce the answer according to its memorized knowledge. To implement the vanilla RAG model, we follow previous work (Lin et al., 2023), use retrieved passages as context and leverage the in-context learning method to conduct the RAG modeling. REPLUG (Shi et al., 2023) is also compared, which ensembles output probabilities from different passage channels. Besides, RA-DIT (Lin et al., 2023) is also compared in our experiments, which optimizes the RAG system using the instruct-tuning method. In our experiments, we reimplement REPLUG and RA-DIT baselines and do not finetune the retriever during our reproduction process, as the retriever we used has already been trained with massive supervised data and is sufficiently strong.
Implementation Details. In our experiments, we employ Minicpm-2.4B-sft (Hu et al., 2024) and lama3-8B-Instruct (Touvron et al., 2023) as backbone models to construct the generation modules, and employ Llama3-8B-Instruct (Touvron et al., 2023) to build the knowledge refinement module. While training the RAG modules with DDR, we use automatic metrics such as Rouge-L and Accuracy to calculate the reward and set $\beta=0.1$. The learning rate is set to 5e-5, and each model is trained for one epoch. For the generation module, we feed 5 retrieved passages as external knowledge for augmenting the generation process. To optimize both knowledge refinement module and generation module, we use LoRA (Hu et al., 2022) for efficient training.
@@ 5 Evaluation Results
In this section, we first evaluate the performance of different RAG methods and then conduct ablation studies to show the effectiveness of different training strategies. Next, we examine the effectiveness of DDR training strategies on the generation module ($V_{\text{Gen}}$) and explore how it balances internal and external knowledge through DDR. Finally, we present several case studies.
@@ 5.1 Overall Performance
Table 2: Ablation Study. Both Vanilla RAG and RAG w/ $V_{\text{KR}}$ are evaluated in a zero-shot setting without any fine-tuning. We then use DDR to optimize the knowledge refinement module ($V_{\text{KR}}$), the generation module ($V_{\text{Gen}}$), and both modules, resulting in three models: RAG-DDR (Only $V_{\text{KR}}$), RAG-DDR (Only $V_{\text{Gen}}$) and RAG-DDR (All).
@@
Vanilla RAG 42.1 78.0 16.6 24.9 22.0 15.1
w/ $V_{\text{KR}}$ 42.2 79.5 16.7 26.7 22.1 14.4
RAG-DDR (Only $V_{\text{KR}}$) 42.5 79.6 16.8 27.3 21.9 15.8
RAG-DDR (Only $V_{\text{Gen}}$) 46.8 81.7 28.3 31.2 32.2 17.0
RAG-DDR (All) 47.0 82.7 28.1 32.5 32.6 17.2
@@
Vanilla RAG 45.4 83.2 20.8 28.5 26.6 12.3
w/ $V_{\text{KR}}$ 46.2 84.0 20.6 30.1 26.9 12.5
RAG-DDR (Only $V_{\text{KR}}$) 46.8 84.7 20.7 30.7 28.5 12.5
RAG-DDR (Only $V_{\text{Gen}}$) 50.2 87.8 25.2 36.9 36.2 14.8
RAG-DDR (All) 50.7 88.2 25.1 37.3 37.3 14.9
@@ 5.2 Ablation Studies
This experiment compares five models, utilizing MiniCPM-2.4b and Llama3-8b to construct the generation module. The Vanilla RAG model relies solely on the generation module ($V_{\text{Gen}}$) to produce answers based on the query and retrieved documents. RAG w/ $V_{\text{KR}}$ adds an additional knowledge refinement module ($V_{\text{KR}}$) to filter the retrieved documents and then feeds query and filtered documents to the generation module ($V_{\text{Gen}}$). RAG-DDR (Only $V_{\text{KR}}$) indicates that we tune the RAG w/ $V_{\text{KR}}$ model using DDR by only optimizing the knowledge refinement module ($V_{\text{KR}}$). RAG-DDR (Only $V_{\text{Gen}}$) only optimizes the generation module ($V_{\text{Gen}}$). RAG-DDR (All) optimizes both $V_{\text{KR}}$ and $V_{\text{Gen}}$.
Compared with the Vanilla RAG model, RAG w/ $V_{\text{KR}}$ improves the RAG performance on almost all evaluation tasks, demonstrating the effectiveness of the knowledge refinement module in improving the accuracy of LLM responses. In contrast, RAG-DDR (Only $V_{\text{Gen}}$) shows greater improvements over RAG w/ $V_{\text{KR}}$ than DDR (Only $V_{\text{KR}}$), indicating that the primary effectiveness of RAG-DDR comes from optimizing the generation module ($V_{\text{Gen}}$) through DDR. When we begin with the RAG-DDR (Only $V_{\text{Gen}}$) model and subsequently optimize the knowledge refinement module, the performance of RAG with $V_{\text{KR}}$ improves. It shows that filtering noise from retrieved documents using feedback from the generation module is effective, which is also observed in previous work (Yu et al., 2023b; Izacard & Grave, 2020). However, the improvements from optimizing the knowledge refinement modules are limited, highlighting that enhancing the generation module’s ability to leverage external knowledge is more critical for the existing RAG system.
@@ 5.3 Characteristics of the Generation Module in RAG-DDR
In this experiment, we explore the characteristics of the generation module ($V_{\text{Gen}}$) by employing various training strategies, including zero-shot (Vanilla RAG), the SFT method (RA-DIT), and DDR (RAG-DDR). As illustrated in Figure 2, we present the performance of $V_{\text{Gen}}$ w/o RAG and $V_{\text{Gen}}$ w/ RAG. These experiments evaluate $V_{\text{Gen}}$’s ability to memorize knowledge and utilize external knowledge. Additionally, we report the average length of responses generated by $V_{\text{Gen}}$.
@@
As shown in Figure 2(a), we compare the performance of the generation module that relies solely on internal knowledge of parametric memory. Compared to the Vanilla RAG model, RA-DIT demonstrates a decline in performance on the NQ and HotpotQA tasks. Such a phenomenon reveals that the model loses previously acquired knowledge while learning new information during SFT (Luo et al., 2023b). In contrast, DDR not only outperforms the RA-DIT method but also achieves consistent improvements over the Vanilla RAG model across all evaluation tasks. This indicates that DDR can help learn more factual knowledge during training while also preventing the loss of previously memorized information through a reinforcement learning-based training approach. Then we feed retrieved documents to the generation module and show the generation performance in Figure 2(b). The evaluation results indicate that RA-DIT marginally outperforms the Vanilla RAG model, while RAG-DDR significantly improves generation accuracy by utilizing factual knowledge from the retrieved documents. Additional experiments showing the general capabilities of RAG-DDR are presented in Appendix A.4.
Finally, we show the average length of responses generated by $V_{\text{Gen}}$ in Figure 2(c). Compared to the Vanilla RAG model, the average length of responses generated by RA-DIT decreases significantly, indicating that the SFT training method tends to cause LLMs to overfit the supervised data. On the contrary, RAG-DDR shows a more similar length distribution with Vanilla RAG model, enabling the model to generate responses of a more appropriate length. It demonstrates that training LLMs to learn data preferences from generated responses can help align the output format of RAG models more closely with that of the original LLMs.
@@
Figure 3: Effectiveness of Different RAG Models in Defending Noisy Information. We use MiniCPM-2.4B to build the generation module ($V_{\text{Gen}}$). And then we retain one informative passage and randomly replace $n$ top-retrieved documents with noisy ones.
@@ 5.4 Effectiveness of RAG-DDR in Using External Knowledge
As shown Table 3, we first show the effectiveness of $V_{\text{Gen}}$ in balancing internal and external knowledge during producing responses. We compare three training strategies: zero-shot (Vanilla RAG), SFT (RA-DIT), and DDR (RAG-DDR). Our experiment establishes three testing scenarios to evaluate the effectiveness of different RAG models by categorizing the evaluation data into three distinct scenarios: Has-Answer, Miss-Answer, and Internal Knowledge. The Has-Answer scenario indicates that the retrieved documents contain the golden answer, which can help the generation module to answer the question accurately. The Miss-Answer scenario indicates that the retrieved documents do not contain the golden answer and fail to provide sufficient support for LLMs to generate accurate responses. Lastly, the Internal Knowledge scenario further evaluates the ability of LLMs in dealing with the conflict between internal and external knowledge. The test cases of the Internal Knowledge scenario indicate that LLMs can generate accurate answers using only parametric memory, whereas RAG models produce incorrect responses.
@@
In the Has-Answer scenario, RAG-DDR and RA-DIT outperform the Vanilla RAG model on all datasets. This indicates that training the generation module enables it to capture essential knowledge facts, improving the accuracy of the generated responses. Compared with RA-DIT, RAG-DDR achieves consistent improvements over the Vanilla RAG model, which shows that DDR can better generalize the external knowledge usage ability to different knowledge-intensive tasks. In the Miss-Answer scenario, all RAG models perform significantly worse compared to the Has-Answer scenario, showing that the retrieved knowledge fails to provide sufficient information for generating accurate responses. However, RAG-DDR and RA-DIT mitigate the performance drop caused by incorporating retrieved documents by fine-tuning the generation module, illustrating their ability to effectively leverage internal knowledge and avoid being misled by irrelevant information. For the Internal Knowledge scenario, we evaluate the ability of RAG models in handling the conflict from external knowledge. The Vanilla RAG model decreases the generation accuracy more than 20% on all tasks, showing that the knowledge conflict can significantly affect the generation result. DDR exhibits strong effectiveness in mitigating the knowledge conflict in the Vanilla RAG model, resulting in a reduction of more than 10% in performance drop. This indicates that DDR effectively balances the utilization of internal and external knowledge, facilitating the robustness of the RAG model.
@@
In the second experiment, we extend the Has-Answer setting and further investigate the denoising effectiveness of $V_{\text{Gen}}$ by adding different number of noise documents. As shown in Figure 3, we increase the noise by randomly replacing $n$ documents from the top-5 retrieved set with the last $n$ documents from the top-100 retrieved candidates, while ensuring that the ground truth document remains in the set. RA-DIT exhibits inconsistent performance and degrades in some RAG scenarios. In contrast, RAG-DDR consistently outperforms the Vanilla RAG model, maintaining a consistent improvement even as the number of noisy documents increases. It further confirms the effectiveness of our RAG-DDR approach in defencing noise.
@@ 5.5 Case Studies
In Table 4, we present three randomly selected examples from the NQ, HotpotQA, and TrivialQA datasets to show the generated responses and evaluate the effectiveness of the RAG-DDR model.
For the first case, the query asks about the “daylight savings time of Canada” and the retrieved documents contain detailed information about when the daylight savings time begins each year in Canada. However, since the exact date of daylight savings time changes annually, the most accurate answer is “the second Sunday in March”. RAG-DDR shows its effectiveness in accurately answering the question, while both Vanilla RAG and RAG-DIT are misled by noisy information such as “14th March” and “10th March”, leading to incorrect responses. This demonstrates that RAG-DDR model has the ability to distinguish the most accurate knowledge from ambiguous or misleading information in the retrieved documents. In the second case, the model must integrate multiple pieces of knowledge from the provided documents to answer the question. While the Vanilla RAG model and RAG-DIT only correctly answer half of the questions using partial knowledge, RAG-DDR successfully identifies the correct start time and end time. This indicates that RAG-DDR has a stronger capacity to integrate factual knowledge from different document segments. As shown in the third case, Vanilla LLM can answer the question correctly only depending on the parametric memory. Nevertheless, both Vanilla RAG and RA-DIT are misled by the confusing information from these retrieved documents and generate the response “the Conservative Party Conference”, which is entirely unrelated to the given question. In contrast, RAG-DDR accurately follows the intent of the question for generating the response, demonstrating the ability of RAG-DDR to mitigate the negative influence of external knowledge.

### v2

@@ 4 Experimental Methodology
This section first describes datasets, evaluation metrics, and baselines. Then, we introduce the implementation details of our experiments. More experimental details are shown in Appendix A.1.
@@
Dataset. In our experiments, we follow RA-DIT (Lin et al., 2023) and use the instruction tuning datasets for training and evaluating RAG models. For all datasets and all baselines, we use bge-large (Xiao et al., 2023) to retrieve documents from the MS MARCO 2.1 (Bajaj et al., 2016).
@@
During the training of DDR, we collect ten datasets covering two tasks, open-domain QA and reasoning. Specifically, we randomly sample 32,805 samples for the training set and 2,000 samples for the development set in our experiments. Following previous work (Lin et al., 2023; Xu et al., 2024a), we select the knowledge-intensive tasks for evaluation, including open-domain question answering, multi-hop question answering, slot filling, and dialogue tasks. The open-domain QA tasks consist of NQ (Kwiatkowski et al., 2019), MARCO QA (Bajaj et al., 2016) and TriviaQA (Joshi et al., 2017), which require models to retrieve factual knowledge to help LLMs answer the given question. For more complex tasks, such as multi-hop QA and dialogue, we use HotpotQA dataset (Yang et al., 2018) and Wikipedia of Wizard (WoW) (Dinan et al., 2019) for evaluation. Besides, we also employ T-REx (Elsahar et al., 2018) to measure one-hop fact look-up abilities of models.
@@
Evaluation. Following Xu et al. (2024a), we utilize Rouge-L and F1 as evaluation metrics for the MARCO QA task and WoW task, respectively. For the rest tasks, we use Accuracy.
@@
Baselines. In our experiments, we compare DDR with five baseline models, including zero-shot models and supervised finetuning models. We first treat the LLM as a black box and conduct three baselines, including LLM w/o RAG, Vanilla RAG and REPLUG. For the LLM w/o RAG model, we directly feed the query to the LLM and ask it to produce the answer according to its memorized knowledge. To implement the vanilla RAG model, we follow previous work (Lin et al., 2023), use retrieved documents as context and leverage the in-context learning method to conduct the RAG modeling. REPLUG (Shi et al., 2023) is compared, which ensembles output probabilities from different passage channels. Self-RAG (Asai et al., 2023) is employed as a baseline, which trains the LLMs to retrieve documents on-demand and reflect on the retrieved documents, and generate responses using a reflection token. RA-DIT (Lin et al., 2023) is also compared in our experiments, which optimizes the RAG system using the instruct-tuning method. In our experiments, we reimplement REPLUG and RA-DIT baselines and don’t finetune the retriever during our reproduction process, as the retriever we use has already been trained with massive supervised data and is sufficiently strong.
Implementation Details. In our experiments, we employ Minicpm-2.4B-sft (Hu et al., 2024) and Llama3-8B-Instruct (Touvron et al., 2023) as backbone models to construct the generation modules and employ Llama3-8B-Instruct (Touvron et al., 2023) to build the knowledge refinement module. More training implementation details of RAG-DDR are presented in Appendix A.1.
@@ 5 Evaluation Results
In this section, we first evaluate the performance of different RAG methods and then conduct ablation studies to show the effectiveness of different training strategies. Then, we examine the effectiveness of DDR training strategies on the generation module ($V_{\text{Gen}}$) and explore how it balances internal and external knowledge through DDR. Finally, we present several case studies.
@@ 5.1 Overall Performance
Table 2: Ablation Study. Both Vanilla RAG (w/o $V_{\text{KR}}$) and Vanilla RAG are evaluated in a zero-shot setting without any fine-tuning. We then use DDR to optimize the knowledge refinement module ($V_{\text{KR}}$), the generation module ($V_{\text{Gen}}$), and both modules, resulting in three models: RAG-DDR (Only Training $V_{\text{KR}}$), RAG-DDR (Only Training $V_{\text{Gen}}$) and RAG-DDR (All Training).
@@
Vanilla RAG (w/o $V_{\text{KR}}$) 42.1 78.0 16.6 24.9 22.0 15.1
Vanilla RAG 42.2 79.5 16.7 26.7 22.1 14.4
RAG-DDR (Only Training $V_{\text{KR}}$) 42.5 79.6 16.8 27.3 21.9 15.8
RAG-DDR (Only Training $V_{\text{Gen}}$) 46.8 81.7 28.3 31.2 32.2 17.0
RAG-DDR (All Training) 47.0 82.7 28.1 32.5 32.6 17.2
@@
Vanilla RAG (w/o $V_{\text{KR}}$) 45.4 83.2 20.8 28.5 26.6 12.3
Vanilla RAG 46.2 84.0 20.6 30.1 26.9 12.5
RAG-DDR (Only Training $V_{\text{KR}}$) 46.8 84.7 20.7 30.7 28.5 12.5
RAG-DDR (Only Training $V_{\text{Gen}}$) 50.2 87.8 25.2 36.9 36.2 14.8
RAG-DDR (All Training) 50.7 88.2 25.1 37.3 37.3 14.9
@@ 5.2 Ablation Studies
This experiment compares five models, utilizing MiniCPM-2.4B and Llama3-8B to construct the generation module. The Vanilla RAG (w/o $V_{\text{KR}}$) relies solely on the generation module ($V_{\text{Gen}}$) to produce answers based on the query and retrieved documents. The Vanilla RAG adds a knowledge refinement module ($V_{\text{KR}}$) to filter the retrieved documents and then feeds query and filtered documents to the generation module ($V_{\text{Gen}}$). RAG-DDR (Only Training $V_{\text{KR}}$) indicates that we tune the Vanilla RAG model using DDR by only optimizing the knowledge refinement module ($V_{\text{KR}}$). RAG-DDR (Only Training $V_{\text{Gen}}$) only optimizes the generation module ($V_{\text{Gen}}$). RAG-DDR (All Training) optimizes both $V_{\text{KR}}$ and $V_{\text{Gen}}$ modules.
Compared with the Vanilla RAG (w/o $V_{\text{KR}}$), Vanilla RAG improves the performance on almost all evaluation tasks, demonstrating the effectiveness of the knowledge refinement module in improving the accuracy of LLM responses. In contrast, RAG-DDR (Only Training $V_{\text{Gen}}$) shows greater improvements over than Vanilla RAG, indicating that the primary effectiveness of RAG-DDR comes from optimizing the generation module ($V_{\text{Gen}}$) through DDR. When we begin with the RAG-DDR (Only Training $V_{\text{Gen}}$) model and subsequently optimize the knowledge refinement module, the performance is further improved. It shows that filtering noise from retrieved documents using feedback from the generation module is effective, which is also observed in previous work (Yu et al., 2023b; Izacard & Grave, 2020). However, the improvements from optimizing the knowledge refinement modules are limited, highlighting that enhancing the generation module’s ability to leverage external knowledge is more critical for the existing RAG system.
@@ 5.3 Characteristics of the Generation Module in RAG-DDR
In this experiment, we explore the characteristics of the generation module ($V_{\text{Gen}}$) by employing various training strategies, including zero-shot (Vanilla RAG), the SFT method (RA-DIT), and DDR (RAG-DDR). As illustrated in Figure 2, we present the performance of $V_{\text{Gen}}$ w/o RAG and $V_{\text{Gen}}$ w/ RAG. These experiments evaluate $V_{\text{Gen}}$’s ability to memorize knowledge and utilize external knowledge. Additionally, we report the average length of responses generated by $V_{\text{Gen}}$ module.
@@
As shown in Figure 2(a), we compare the performance of the generation module that relies solely on internal knowledge of parametric memory. Compared to the Vanilla RAG model, RA-DIT demonstrates a decline in performance on the NQ and HotpotQA tasks. Such a phenomenon reveals that the model loses previously acquired knowledge while learning new information during SFT (Luo et al., 2023b). In contrast, DDR not only outperforms the RA-DIT method but also achieves consistent improvements over the Vanilla RAG model across all evaluation tasks. This indicates that DDR can help LLMs learn more factual knowledge during training while also preventing the loss of previously memorized information through a reinforcement learning-based training approach. Then we feed retrieved documents to the generation module and show the generation performance in Figure 2(b). The evaluation results indicate that RA-DIT marginally outperforms the Vanilla RAG model, while RAG-DDR significantly improves generation accuracy by utilizing factual knowledge from the retrieved documents. Additional experiments showing the general capabilities of RAG-DDR are presented in Appendix A.7.
Finally, we show the average length of responses generated by $V_{\text{Gen}}$ in Figure 2(c). Compared to the Vanilla RAG model, the average length of responses generated by RA-DIT decreases significantly, indicating that the SFT training method tends to cause LLMs to overfit the training dataset. On the contrary, RAG-DDR shows a more similar length distribution with the Vanilla RAG model, enabling the model to generate responses of a more appropriate length. It demonstrates that training LLMs to learn data preferences from generated responses can help align the output format of RAG models more closely with that of the original LLMs.
@@
Figure 3: Effectiveness of Different RAG Models in Defending Noisy Information. We use MiniCPM-2.4B to build the generation module ($V_{\text{Gen}}$). Then we retain one informative document and randomly replace $n$ top-retrieved documents with noisy ones.
@@ 5.4 Effectiveness of RAG-DDR in Using External Knowledge
As shown in Table 3, we first show the effectiveness of $V_{\text{Gen}}$ in balancing internal and external knowledge during producing responses. We compare three training strategies: zero-shot (Vanilla RAG), SFT (RA-DIT), and DDR (RAG-DDR). Our experiment establishes three testing scenarios to evaluate the effectiveness of different RAG models by categorizing the evaluation data into three distinct scenarios: Has-Answer, Miss-Answer, and Internal Knowledge. The Has-Answer scenario indicates that the retrieved documents contain the golden answer, which can help the generation module to answer the question accurately. The Miss-Answer scenario indicates that the retrieved documents do not contain the golden answer and fail to provide sufficient support for LLMs to generate accurate responses. Lastly, the Internal Knowledge scenario further evaluates the ability of LLMs in dealing with the conflict between internal and external knowledge. The test cases of the Internal Knowledge scenario indicate that LLMs can generate accurate answers using only parametric memory, whereas RAG models may produce incorrect responses.
@@
In the Has-Answer scenario, RAG-DDR and RA-DIT outperform the Vanilla RAG model on all datasets. This indicates that training the generation module enables it to capture essential knowledge facts, improving the accuracy of the generated responses. Compared with RA-DIT, RAG-DDR achieves consistent improvements over the Vanilla RAG model, which shows that DDR can better generalize the external knowledge usage ability to different tasks. In the Miss-Answer scenario, all RAG models perform significantly worse compared to the Has-Answer scenario, showing that the retrieved knowledge fails to provide sufficient information for generating accurate responses. Compared to the Vanilla RAG, RAG-DDR and RA-DIT mitigate the performance drop caused by incorporating retrieved documents by fine-tuning the generation module, illustrating their ability to effectively leverage internal knowledge and avoid being misled by irrelevant information. For the Internal Knowledge scenario, we evaluate the ability of RAG models in handling the conflict from external knowledge. The Vanilla RAG model decreases the generation accuracy more than 20% on all tasks, showing that the knowledge conflict can significantly affect the generation result. DDR exhibits strong effectiveness in mitigating the knowledge conflict in the Vanilla RAG model, resulting in a reduction of more than 10% in performance drop. This indicates that DDR effectively balances the internal and external knowledge, facilitating the robustness of the RAG model.
@@
In the second experiment, we extend the Has-Answer setting and further investigate the denoising effectiveness of $V_{\text{Gen}}$ by adding the different number of noise documents. As shown in Figure 3, we increase the noise by randomly replacing $n$ documents from the top-5 retrieved set with the last $n$ documents from the top-100 retrieved candidates, while ensuring that the ground truth document remains in the set. RA-DIT exhibits inconsistent performance and degrades in some RAG scenarios. In contrast, RAG-DDR consistently outperforms the Vanilla RAG model, maintaining a consistent improvement even as the number of noisy documents increases. It further confirms the effectiveness of our RAG-DDR approach in defending noise.
@@ 5.5 Case Studies
In Table 4, we present three randomly selected examples from the NQ, HotpotQA, and TriviaQA datasets to show the generated responses and evaluate the effectiveness of the RAG-DDR model.
For the first case, the query asks about the “daylight savings time of Canada” and the retrieved documents contain detailed information about when the daylight savings time begins each year in Canada. However, since the exact date of daylight savings time changes annually, the most accurate answer is “the second Sunday in March”. RAG-DDR shows its effectiveness in accurately answering the question, while both Vanilla RAG and RA-DIT are misled by noisy information such as “14th March” and “10th March”, leading to incorrect responses. This demonstrates that RAG-DDR model has the ability to distinguish the most accurate knowledge from ambiguous or misleading information in the retrieved documents. In the second case, the model must integrate multiple pieces of knowledge from the provided documents to answer the question. While the Vanilla RAG model and RAG-DIT only correctly answer half of the questions using partial knowledge, RAG-DDR successfully identifies the correct start time and end time. This indicates that RAG-DDR has a stronger capacity to integrate factual knowledge from different document segments. As shown in the third case, Vanilla LLM can answer the question correctly only depending on the parametric memory. Nevertheless, both Vanilla RAG and RA-DIT are misled by the confusing information from these retrieved documents and generate the response “the Conservative Party Conference”, which is entirely unrelated to the given question. In contrast, RAG-DDR accurately follows the intent of the question for generating the response, demonstrating the ability of RAG-DDR to mitigate the negative influence of external knowledge.